OpenStudy (anonymous):

Find f'(x) if f(x) = ln((4x+9)/(5x-7))^(1/2)

OpenStudy (anonymous):

Do you do U' over U?

OpenStudy (anonymous):

Anyone?

zepdrix (zepdrix):

\[\frac{ \ln (4x+9) }{ \sqrt{5x-7} }\] This is the problem? :D

OpenStudy (anonymous):

ln of (4x+9)/(5x-7) all square rooted :D

OpenStudy (anonymous):

So the 4x+9 is raised to the power of 1/2 as well.

zepdrix (zepdrix):

kk :) So your OUTERMOST function is Natural Log, so You'll take the derivative of the natural log, then we'll have a nasty chain rule inside of there ^^

zepdrix (zepdrix):